OpenStudy (anonymous):

A box with no top is made by taking a rectangular sheet of metal and cutting away four equal-sized squares (one from each corner of the sheet) and then folding up the resulting aps. The box should be made 4 cm deep and the length of the base should be 15 cm more than the width. If the volume is to be 1000 then what should be the width of the sheet of metal?

OpenStudy (anonymous):

ok so volume here is 1000. 1000=length*width*height, say width=x then length=x+15. so solve the quadratic equation 4x^2+20x-1000=0

OpenStudy (anonymous):

i got 26?

OpenStudy (anonymous):

well that wont give you the right volume tho

OpenStudy (anonymous):

actually nevermind the quadratic formula , i think you need help its late at night im tired, there answer is 10. 10*25*4 is 1000

OpenStudy (anonymous):

10 is the width length is 15 more so its 25. cheers

OpenStudy (anonymous):

thats not one of the mutiple choices... but thanks for trying

OpenStudy (anonymous):

lol what are your multiple choices

OpenStudy (anonymous):

list them

OpenStudy (anonymous):

Still working on this one?

OpenStudy (anonymous):

yes he wants to know the width still, i said it was 10 but i guess that is wrong

OpenStudy (anonymous):

A) 18 cm B) 20 cm C) 22 cm D) 24 cm E) 26 cm

OpenStudy (anonymous):

oh i forgot about the boxes that were cut out, yes 10 is the width of the inner box but its made from the sheet metal so you have to add the little boxes you cut off which makes it 10+4+4=18, however my reasoning was correct to get 10. i just failed to realize that thats not what the question asked for.

OpenStudy (anonymous):

ok, well thanks...

OpenStudy (anonymous):

do you understand better how to do this problem now tho?

OpenStudy (anonymous):

not at all